OpenStudy (anonymous):

5 20 -- div -- divide and simplify s s

OpenStudy (anonymous):

panther please try it by urself..

OpenStudy (dumbcow):

Flip and Multiply

OpenStudy (anonymous):

division of fractions means the multiplication with reciprocals...now go ahead.

OpenStudy (anonymous):

u try it, we are here to help u..but u first try it.

OpenStudy (anonymous):

hello @panther...

OpenStudy (anonymous):

so i flip both and multiply

OpenStudy (anonymous):

no flip only the one which is after division sign

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

thats my answer

OpenStudy (anonymous):

\[\frac5s \div\frac{20}s\]

OpenStudy (anonymous):

yes u are correct.

OpenStudy (anonymous):

\[\frac5s \div \frac{20}s =\frac5s * \frac s{20}=\frac14\]

OpenStudy (anonymous):

see u have done it by urself..

OpenStudy (anonymous):

which is why i do like it when i see how its done as opposed to getting the answer

OpenStudy (anonymous):

can you help with my word problem